# Extreme Points and Large Contests*

### v1

@@ Abstract.
*I am particularly indebted to Ioan Manolescu for his generous time. I also thank Yaron Azrieli, Christian Chambers, Christian Ewerhart, Byeong-hyeon Jeong, Patrick Lahr, Igor Letina, Brendan Pass, Philipp Strack and Stefan Wenger for helpful discussions.
@@ 2. Extreme Points
$\mathcal{F}^{*}=\left\{f\in\mathcal{F}\ :\int_{\mathcal{X}}f(x,\mu)\ d\mu(x)\leq k,\ k\in(0,1)\right\}\ .$ (1)
@@
Let $\Phi:\mathcal{X}\times\Delta(\mathcal{X})\longrightarrow[0,1]$. Then,
@@ 3.1. Complete Information
Denote $\mathbf{M}$ for a probability measure on $(\Delta(\mathcal{X}),\mathcal{B}(\Delta(\mathcal{X})))$. [^9] Endowed with the topology of weak convergence, $\Delta(\mathcal{X})$ has a natural notion of $\sigma$-algebra. Formally, $\mu_{n}\rightarrow\mu\iff\int_{\mathcal{X}}gd\mu_{n}\rightarrow\int_{\mathcal{X}}gd\mu\ ,\ g\ \text{continuous}$. The payoff of the designer is given by
@@
Proof.
In order to apply Theorem 2, we need to show that (5) is convex in $q$ and supermodular in $(q,x)$. We have the following result.
Lemma 1.
$x^{*}(q)$ is non-decreasing. In particular, $dx^{*}(q)/dq=0$.
Proof.
The form of $q^{*}$ directly follows from Theorems 1 and 2. We only need to show the following result.
Lemma 2.
$q^{*}$ is independent of $\mathbf{M}$.
Proof.
Proposition 1 characterizes the optimal allocation rules. It provides an economic foundation for the allocation rules discussed in the Introduction. Moreover, Proposition 1 significantly simplifies the characterization of equilibria in this class of games with complete information. To this end, consider the case in which $\pi(x)$ is increasing. Then, it is straightforward to derive the following result.
@@
Similarly, when $\pi(x)$ is non-monotone, the designer might prefer agents to exert some intermediate level of effort. Then, by setting the two thresholds appropriately, the designer can induce agents to do so by extending the jump in the middle, so as not to give them the incentive to slightly increase their effort level at any given $\mu$ they face.
@@ 3.2. Incomplete Information
Proposition 1 still holds under incomplete information.
@@ 3.3. More General Objective Function
In other words, the designer might be interested in maximizing some aggregate measure of effort, and s/he might also be concerned with the agents who do not win a prize. The result derived in Proposition 1 goes through also for the specification in (10), both when agents utilities are given by (4) and (9).
@@
The Allocation Rule which maximizes (10) subject to (3) is the same as in Proposition 1.
@@ A.1. Proof of Theorem 1
Several steps are involved.
@@
$\mathcal{F}$ is a vector space.
Define
$\mathcal{H}=\left\{f:\mathcal{X}\times\mathcal{M}_{\text{signed}}\longrightarrow\mathbb{R}\ ,\ f\ \text{bounded}\right\}\ ,$ (11)
where $\mathcal{M}_{\text{signed}}$ is the set of finite signed measures. We give it the linear structure. Then, since $\Delta(\mathcal{X})$ is a subset of $\mathcal{M}_{\text{signed}}$, it follows that $\mathcal{F}\subseteq\mathcal{H}$. Hence, $\mathcal{F}$ is a vector space.
Step 2.
$\mathcal{F}$ is compact and convex.
Lemma 4.
$\mathcal{F}$ is compact with respect to the weak topology.
Proof.
Define $\mathcal{F^{\prime}}=\left\{f:\mathcal{X}\longrightarrow[0,1],f\ \text{increasing}\right\}$. $\mathcal{F}$ can be rewritten as
$\mathcal{F}$ $=\{\psi:\underbrace{\Delta(\mathcal{X})}_{B}\rightarrow\underbrace{\mathcal{F^{\prime}}}_{A}\}$ (12)
$=A^{B}$ (13)
$\mathcal{F^{\prime}}$ endowed with the $L^{1}$ norm is a Banach space and it is compact with respect to the topology induced by this norm. Since $\mathcal{X}$ is finite dimensional, this topology is equivalent to the weak topology.
Since $\mathcal{M}_{\text{signed}}$ is a Banach space with respect to the total variation norm, $\Delta(\mathcal{X})$ is also a Banach space. Moreover, it is compact with respect to the weak topology.
Tychonoff’s Theorem implies that $\mathcal{F}$ is compact in the product topology of weak topologies, since it is a product of sets which are compact in the weak topology. However, by Theorem 3.10 in Brezis (2011), the product topology of weak topologies is the weak topology of the product space. This shows the compactness of $\mathcal{F}$. ∎
Step 3.
@@
A theorem in Dubins (1962) shows that an extreme point of $\mathcal{F}^{*}$ is a convex combination of at most two extreme points of $\mathcal{F}$. Hence, extreme points of $\mathcal{F}^{*}$ are of the form
@@
We provide an alternative proof of Proposition 1, without using Theorem 2. By Theorem 1, we know that $\mathcal{F}$ is a compact and convex. It is straightforward to show that $\mathcal{F}^{*}$ is also compact and convex. Convexity follows form the linearity of the constraints. Moreover, since $\mathcal{F}^{*}$ is closed and $\mathcal{F}^{*}\subseteq\mathcal{F}$, then $\mathcal{F}^{*}$ is compact.
@@
Theorem 0 (Bauer’s Maximum Principle).
Let $A$ be a locally convex, topological vector space, $B$ be a non-empty compact, convex subset of $A$, and $S:B\longrightarrow\mathbb{R}$ be a continuous linear function. Then the set $E$ of maximizers of $S$ over $B$ is a face of $B$. Furthermore, $E$ contains an extreme point of $B$.
Therefore, we know that maximizers of (2) are of the form of Theorem 1. In order to characterize more precisely the maximizers depending on whether (2) is increasing or decreasing, we first derive the following useful result.
@@
$=\lambda\Pi(f_{1})+(1-\lambda)\Pi(f_{2})$
$=\lambda\Pi(k_{1})+(1-\lambda)\Pi(k_{2})\ .$
We now show that the payoff $\Pi(f^{*})$ is maximal.

### v2

@@ Abstract.
*I am particularly indebted to Ioan Manolescu for his generous time. I also thank Yaron Azrieli, Christian Chambers, Christian Ewerhart, Dominik Inauen, Byeong-hyeon Jeong, Patrick Lahr, Igor Letina, Brendan Pass, Philipp Strack and Stefan Wenger for helpful discussions.
@@ 2. Extreme Points
$\mathcal{F}^{*}=\left\{f\in\mathcal{F}\ :\int_{\mathcal{X}}f(x,\mu)\ d\mu(x)\leq k,\ k\in(0,1)\right\}\ .$
@@
Let $\Phi:\mathcal{X}\times\Delta(\mathcal{X})\times\mathcal{X}\longrightarrow[0,1]$. Then,
@@ 3.1. Complete Information
Denote $\mathbf{M}$ for a probability measure on $(\Delta(\mathcal{X}),\mathcal{B}(\Delta(\mathcal{X})))$. [^9] Endowed with the topology of weak convergence, $\Delta(\mathcal{X})$ has a natural notion of $\sigma$-algebra. Formally, $\mu_{n}\rightarrow\mu\iff\int_{\mathcal{X}}gd\mu_{n}\rightarrow\int_{\mathcal{X}}gd\mu\$, for any $g$ continuous. The payoff of the designer is given by
@@
Theorem 3 characterizes the optimal allocation rules. It provides an economic foundation for the allocation rules discussed in the Introduction. Moreover, Theorem 3 significantly simplifies the characterization of equilibria in this class of games with complete information. To this end, consider the case in which $\pi(x)$ is increasing. Then, it is straightforward to derive the following result.
@@
Similarly, when $\pi(x)$ is non-monotone, the designer might prefer agents to exert some intermediate level of effort. Then, by setting the two thresholds appropriately, the designer can induce agents to do so by extending the jump in the middle, so as not to give them the incentive to slightly increase their effort level, for any given competition $\mu$ they face.
@@ 3.2. Incomplete Information
Theorem 3 still holds under incomplete information.
@@ 4.1. More General Objective Function
In other words, the designer might be interested in maximizing some aggregate measure of effort, and s/he might also be concerned with the agents who do not win a prize. The result derived in Theorem 3 goes through also for the specification in (9), both when agents utilities are given by (3) and (8).
@@
The Allocation Rule which maximizes (9) subject to (2) is the same as in Theorem 3.
@@
4.2. Multiple Prize Levels
Suppose there are different prize levels $\mathbf{k}=\{k_{1},k_{2},\dots k_{n}\}$, such that $\sum_{i}^{n}k_{i}=k\in(0,1)$, with respective values $V_{1},V_{2},\dots,V_{n}$. The following result shows that it is without loss of generality for the contest designer to focus on a single prize.
Lemma 2.
$k\mapsto\Pi(k)$ is concave.
Proof.
$=\lambda\Pi(f_{1})+(1-\lambda)\Pi(f_{2})$
$=\lambda\Pi(k_{1})+(1-\lambda)\Pi(k_{2})\ .$
4.3. Large Contests with Diversity Constraints
Consider now the following variation of the model. In addition to choosing an effort level $x\in\mathcal{X}$, suppose agents have an observable characteristic $\theta\in\Theta=\{\theta_{1},\theta_{2}\}$, with relative shares $\alpha_{1},\alpha_{2}$, respectively. Assume the contest designer faces an additional constraint to give a share $h$ of the prizes to the subset of agents with characteristic $\alpha_{1}$:
$\int_{\mathcal{X}}q(x,\mu_{|\theta_{1}})\ d\mu_{|\theta_{1}}(x)=h\ ,$
with $k>\alpha_{1}>h$.
This additional constraint might be interpreted as a diversity constraint, or a quota on the number of agents with a given characteristic.
For example, at EPFL in Switzerland, the number of students admitted to the Bachelor’s programmes has recently been capped at 3000: Swiss students holding the relevant qualifications will be admitted without limitations, whereas foreign students will be admitted according to the number of remaining places. [^12] See https://actu.epfl.ch/news/admissions-to-bachelor-programs-will-be-limited-st/.
The payoff of the designer can now be written as
$\Pi$ $=\int_{\Delta(\mathcal{X})}\int_{\mathcal{X}}q(x,\mu)\pi(x)\ d\mu(x)\ d\mathbf{M}(\mu)$
$=\int_{\Delta(\mathcal{X})}\alpha_{1}\int_{\mathcal{X}}q(x,\mu_{|\theta_{1}})\pi(x)\ d\mu_{|\theta_{1}}(x)+\alpha_{2}\int_{\mathcal{X}}q(x,\mu_{|\theta_{2}})\pi(x)\ d\mu_{|\theta_{2}}(x)\ d\mathbf{M}(\mu)$
$=\int_{\Delta(\mathcal{X})}\alpha_{1}\Pi(\mu_{|\theta_{1}})+\alpha_{2}\Pi(\mu_{|\theta_{2}})\ d\mathbf{M}(\mu)\ .$
The designer’s problem is then equivalent to two subproblems, one for each subset of agents $\mu|_{\theta_{i}}$, $i=1,2$.
@@ A.1. Proof of Theorem 1
Define $\mathcal{F^{\prime}}=\left\{f:\mathcal{X}\longrightarrow[0,1],f\ \text{increasing}\right\}$.
@@
A theorem in Winkler (1988) shows that an extreme point of $\mathcal{F}^{*}$ is a convex combination of at most two extreme points of $\mathcal{F}$. Hence, extreme points of $\mathcal{F}^{*}$ are of the form
@@
Several steps are involved.
Step 2.
$\mathcal{F}$ is a subset of a vector space.
Define
$\mathcal{H}=\left\{f:\mathcal{X}\times\mathcal{M}_{\text{signed}}\longrightarrow\mathbb{R}\ ,\ f\ \text{bounded}\right\}\ ,$
where $\mathcal{M}_{\text{signed}}$ is the set of finite signed measures. We give it the linear structure. Then, since $\Delta(\mathcal{X})$ is a subset of $\mathcal{M}_{\text{signed}}$, it follows that $\mathcal{F}\subseteq\mathcal{H}$. Hence, $\mathcal{F}$ is a subset of a vector space.
Step 3.
$\mathcal{F}$ is compact and convex.
Lemma 4.
$\mathcal{F}$ is compact with respect to the product topology.
Proof.
$\mathcal{F}$ can be rewritten as
$\mathcal{F}$ $=\{\psi:\Delta(\mathcal{X})\rightarrow\mathcal{F^{\prime}}\}$
$=\prod_{\mu\in\Delta(\mathcal{X})}\mathcal{F}^{\prime}\ .$
Since each $f\in\mathcal{F}^{\prime}$ is bounded and uniformly integrable, by Dunford-Pettis Theorem $\mathcal{F}^{\prime}$ is compact with respect to the weak topology on $L^{1}(\mu)$, for each $\mu\in\Delta(\mathcal{X})$. Moreover, $\Delta(\mathcal{X})$ is compact with respect to the weak topology.
Tychonoff’s Theorem implies that $\mathcal{F}$ is compact in the product topology of weak topologies. ∎
Step 4.
Bauer’s Maximum Principle.
It is straightforward to show that $\mathcal{F}^{*}$ is also compact and convex. Convexity follows form the linearity of the constraints. Moreover, since $\mathcal{F}^{*}$ is closed and $\mathcal{F}^{*}\subseteq\mathcal{F}$, then $\mathcal{F}^{*}$ is compact.
@@
Theorem 0 (Bauer’s Maximum Principle Ball (2023)).
Let $A$ be a locally convex, topological vector space, $B$ be a non-empty compact, convex subset of $A$, and $S:B\longrightarrow\mathbb{R}$ be a quasiconvex function. Then the set $E$ of maximizers of $S$ over $B$ is a face of $B$. Furthermore, $E$ contains an extreme point of $B$.
Clearly, (1) is continuous with respect to the weak topology on $L^{1}(\mu)$ for each $\mu\in\Delta(\mathcal{X})$; hence, it is continuous with respect to the product topology of weak topologies. Therefore, we know that maximizers of (1) are of the form of Theorem 1.
The following results shows that maximizers of (1) are also maximizers of (4).
@@
Lastly, in order to apply Theorem 2, we need to show that (4) is convex in $q$ and supermodular in $(q,x)$. We have the following result.
Lemma 6.
$x^{*}(q)$ is non-decreasing. In particular, $dx^{*}(q)/dq=0$.
Proof.
A.3. Alternative Proof of Theorem 3.
Proof.
We provide an alternative proof of Theorem 3, without using Theorem 2.